OpenStudy (anonymous):

1. Whats the molality of a solution of 2.0g of NaCl dissolved in 1.0 kg of water? 2. If a sample of gas occupies a volume of 100 mL at a pressure of 200 mm Hg at at constant temperature and the pressure is changed to 400 mm Hg, the volume will change to?

OpenStudy (anonymous):

@eric_d Know these?

OpenStudy (abhisar):

Hello @KissesFromKristine ! \(\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\\\color{white}{.}\\\Huge\sf\color{blue}{~~~~Welcome~to~OpenStudy!~\ddot\smile}\\\color{white}{.}\\\\\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\)

OpenStudy (eric_d):

 what do u think @Abhisar

OpenStudy (abhisar):

\[Molality = \frac{ No. of moles of solute }{ Mass of solvent \in Kg }\]

OpenStudy (abhisar):

So, molality of 2 g NaCl in 1 kg of water will be 0.034

OpenStudy (abhisar):

are u getting guys ?

OpenStudy (eric_d):

how do u et 0.034

OpenStudy (eric_d):

get*

OpenStudy (eric_d):

@Abhisar

OpenStudy (eric_d):

so I got to convert 2g to mole first

OpenStudy (anonymous):

Thanks for the help @Abhisar, I ended up figuring that one out though. :P Know how to do the second one?

OpenStudy (abhisar):

okay... 58.5 g NaCl = 1 mole then 2g will contain how many moles ? @eric_d

OpenStudy (abhisar):

yes @KissesFromKristine for the second one the answer will be 50 ml btw what's the answer for 1st one ?

OpenStudy (abhisar):

For the second one u can use the formula \[\frac{ P1 }{ P2 } = \frac{ V2 }{ V1 }\]

OpenStudy (anonymous):

The first one is what you said: 0.034

OpenStudy (abhisar):

okay...u got the second one ?

OpenStudy (abhisar):

U can derive the formula i said from the equation PV=nRT

OpenStudy (anonymous):

Yep, makes sense when you put up the formula, I forgot about that.

OpenStudy (anonymous):

you're really good at chem, thanks for all the help! :) <3

OpenStudy (eric_d):

I got it @Abhisar Thanks for the explanation

OpenStudy (abhisar):

\(\Huge\text{Anytime !}\) \(\huge\ddot\smile\)